### v1

# Quantum response theory and momentum-space gravity M. Mehraeen

Abstract
We present a quantum response approach to momentum-space gravity in dissipative multiband systems, which dresses both the quantum geometry–through an interband Weyl transformation–and the equations of motion. In addition to clarifying the roles of the contorsion and symplectic terms, we introduce the three-state quantum geometric tensor and discuss the significance of the emergent terms from a gravitational point of view. We also identify a dual quantum geometric drag force in momentum space that provides an entropic source term for the multiband matrix of Einstein field equations.
Suggested keywords
Introduction.The space of physical states of a quantum mechanical system–complex projective Hilbert space–has the structure of a Kahler manifold [1]. This has inspired various geometric and topological interpretations of physical phenomena in quantum systems over the years [2, 3]. And, in several subfields of condensed matter physics, this quantum state geometry has recently garnered immense interest [4, 5, 6, 7, 8], with the identification of the role of the quantum metric [9]–the real part of the quantum geometric tensor (QGT)–in an increasing number of physical effects [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
While the dynamical significance of the Berry curvature has been known for decades [23, 24, 25, 26, 27, 28], it is perhaps the emergence of the quantum-metric Levi-Civita connection [29] that truly reveals the geometric underpinning of carrier dynamics and highlights the geodesic nature of the motion of Bloch electrons. Combining this with the Einstein field equations (EFE) arising from the Riemannian structure of quantum state manifolds has led to the intriguing recent proposal of momentum-space gravity [30] in condensed matter systems. Given the semiclassical nature of this approach based on wavepacket dynamics, a natural question that arises is the form this theory of gravity takes in momentum space from the vantage point of quantum response theory.
Here, we address this question through a diagrammatic approach, with the general idea underlying this work presented schematically in Fig. 1. We first apply Kubo formulas to propose a diagrammatic generalization of carrier dynamics to multiband systems in the presence of finite dissipation. To this end, we build on recent developments [31, 32, 33, 34, 35] in two-state 111Here, we refer to an -state quantum geometric object as one that carries band indices. For example, and are the single- and two-state quantum metric tensors, respectively. quantum geometry to propose the three-state QGT and show that it appears already at the quadratic-response level. We also discuss the appearance of the quantum geometric contorsion tensor [1] and analogous symplectic terms when the full Berry covariant derivative is used to derive the Hermitian connection components.
Physically, much intuitive insight has been gained from viewing the Berry curvature as an intrinsic gauge field strength tensor in momentum space [3]. Following this, and the standard argument in general relativity [37, 38], we argue that once the choice of connection is made to be Levi-Civita, the remaining terms from the Hermitian connection and three-state QGT in the carrier position’s equation of motion can then be interpreted as intrinsic matter fields arising from the multistate quantum geometry.
[…]
Figure 1: Schematic summary of this work. Correlation functions corresponding to conductivity Kubo formulas are applied to simultaneously dress with dissipation the quantum geometry, carrier dynamics and momentum-space field equations. The latter two yield dressed versions of the two central equations of general relativity, resulting in dressed gravity in momentum space.
A noteworthy feature of the present approach is the appearance of dressed quantum geometric quantities in the equations of motion and EFE. While we discuss the gravitational significance of this later on, here we mention that there is increasing ongoing effort to define quantum geometry in interacting and disordered systems [39, 40, 41, 42, 43, 44, 45]. The approach which naturally emerges in the present derivation is closest to that of Ref. [40] and, with the gravitational context in mind, is essentially a dissipation-induced multiband Weyl transformation of the quantum geometry. We apply this result to show that the dissipative correction to the EFE can be related to the entropic cost of dressing the metric through a dual momentum-space drag force induced by the dissipative scattering. This, in turn, can be thought of as providing a diagrammatic generalization of the corresponding semiclassical arguments to dissipative multiband systems. We conclude by presenting an outlook on possible insights that may be obtained from these results.
Multistate quantum geometry.In the presence of an applied electric field, the electric dipole interaction is included to the system Hamiltonian, where is the position operator. In the k-space representation, this takes the form [46]
, where is the Berry connection and the periodic part of the Bloch state 222Regarding notation, we use Greek letters for real- or momentum-space indices–with Einstein summation implied–and Latin letters for band indices. Furthermore, and . We henceforth work locally in momentum space and thus drop the momentum parameter from the notation. The commutator of the position operator and local momentum-space operators naturally defines a Berry covariant derivative
, which takes the form
It should be stressed that this includes the full Berry connection, which is decomposed into diagonal and off-diagonal elements in band space as
The transition dipole matrix element can be viewed as a component of the complex-valued vielbein matrix in band space [31], thereby providing a geometric rationale for using off-diagonal Berry connection components as fundamental elements in the construction of multistate quantum geometric quantities. And the Hilbert-Schmidt inner product
induces complex Riemannian structure on the manifold of quantum states and allows for the definition of quantum geometric invariants. The simplest such two-state object is the well-known QGT, which is obtained from the inner product of tangent basis vectors,
, and which obeys the projector calculus identity
with the projection operator given by
. Note that for , Eq. (2) may be regarded as an alternative definition of the QGT in the projector calculus language. Indeed, this approach can be used to identify other elements of two-state quantum geometry [34, 33], with the two-state QGT being the simplest member. The real and imaginary parts of this tensor define the two-state quantum metric and Berry curvature tensors,
and
. Furthermore, we note that the QGT is basically the band resolution of single-state quantum geometry, and yields the familiar Fubini-Study metric and (single-state) Berry curvature tensors once one sums over intermediate states, as is done in the second term on the right in Eq. (2).
Moving beyond the QGT, the position operator acting on the tangent basis vectors induces a Hermitian connection [31] which parallel-transports the QGT and whose components are given by
. Using the standard definition of the (two-state) torsion tensor,
, we note that the real part of the Hermitian connection is expressed as
where
is a Levi-Civita connection component of the band-resolved quantum metric and the second term, given by
is identified as the quantum geometric contorsion tensor [1], which contains the torsionful part of the metric connection and arises once one uses the full Berry covariant derivative to define the Hermitian connection, as opposed to using only the diagonal Berry connection components. We note that the contorsion tensor inherits the antisymmetric property
from the torsion tensor.
Motivated by a geometric characterization of nonlinear responses, we now present an extension of this framework beyond the two-state formalism by introducing the three-state QGT
which cyclically connects three distinct states, and represents a natural generalization of its two-state counterpart. Formally, this can be defined by utilizing the non-Abelian QGT [9, 48, 32, 35] in operator form, composed of basis vector derivatives, as
, which allows for the definition
with
. This naturally generalizes the two-state projector identity given by Eq. (2) to the three-state identity
lending further support to the proposition that the three-state QGT defined above is indeed a central object in higher-state quantum geometry.
We point out that the real and imaginary parts of the three-state QGT are decomposed into linear combinations of the non-Abelian quantum metric and Berry curvature tensors as
thereby exhibiting a more complex structure compared to the two-state counterpart. The symmetric and antisymmetric parts,
and
, are also expressible in terms of these quantities. Specifically, we note that the antisymmetric term is the band resolution of the torsion tensor, i.e.,
, while
is its symmetric counterpart. As is shown in the next section, the three-state QGT already makes an appearance at the quadratic-response level in dissipative multiband systems and is required for a complete geometric characterization of the response functions.
Dressed carrier dynamics from diagrammatics.Motivated by recent developments via density-matrix methods in generalizing carrier dynamics [49, 50], here, we obtain the equation of motion for the carrier position by taking the route of Kubo formulas. Using recently developed diagrammatic approaches within the Matsubara formalism [51, 52], the linear and quadratic ac conductivities are obtained 333See the Supplemental Material for details on the evaluation of the response functions, equations of motion, renormalization functions and the dissipative EFE.. In order to obtain dissipative carrier dynamics, we note that the dc current density,
, is related to the equation of motion of the carrier position and the carrier density as
[49, 50]. Therefore, upon evaluating the correlation functions, and using the equation of motion of the carrier momentum,
we arrive at the equation of motion for the carrier position, which is expressed as
where is the group velocity and the ’s are various renormalization factors that depend on the dissipation parameter. Eq. (9) is one of the central results of this work and represents a diagrammatic generalization of carrier dynamics to dissipative multiband systems. As can be seen, in addition to dressing the known terms related to the Berry curvature and Levi-Civita connection, which are interpreted as the (multiband) momentum-space magnetic field and geodesic contributions, respectively, several additional contributions are identified. The remainder of this section is devoted to a discussion of these terms and their physical significance, particularly within the context of the gravitational interpretation.
The matrix (with
) couples to the Christoffel symbol in the equation of motion and may thus be regarded as the local multiband generalization of the effective mass term in Ref. [30]. Physically, this implies that the effective gravitational force the electron in band experiences from the Levi-Civita connection between bands and is inversely proportional to the local energy difference between the two bands. Interestingly, this force can be attractive or repulsive depending on the sign of . And since the effective mass is a local quantity, its variation also contributes to the equation of motion, which explains the
term in the equation of motion.
It is worth elaborating further on the appearance of dressed quantum geometric quantities in Eq. (9) as opposed to bare ones. With the exception of the Levi-Civita connection, we define dressed quantum geometric quantities as
where is a dimensionless measure of the dissipation strength, with measuring the self-energy. We stress that dressing the quantum geometry is an essential inclusion for a gravitational interpretation to hold, as, in general relativity, the matter distribution and the (spacetime) curvature are related and influence each other [37]. In our case, the introduction of scatterers to the system has the effect of screening the quantum geometry through the interband scattering function
, which acts as a dissipation-induced multiband Weyl transformation on the various geometric terms in the equation of motion. We also note that this choice of dressing has the favorable property of preserving the symmetries of the quantum metric and Berry curvature in band space.
The additional quantum geometric objects that appear in Eq. (9) are given by
which contains the contribution from the rest of the dressed Hermitian connection, i.e., the contorsion and its symplectic counterpart, and
which encapsulates the contributions from the dressed three-state QGT 444Note that we define the dressed three-state QGT in terms of the first two band indices as . Together, and can be intuitively viewed as describing intrinsic momentum-space matter fields [37, 38] that help capture the quantum geometry of Bloch electrons, including the symplectic and non-Abelian contributions, at second order in the carrier momentum.
Finally, consider now the term proportional to in Eq. (9), which includes a dissipative linear-response addition to the equation of motion
that is also not captured in the semiclassical or density-matrix approaches. Applying the dual-space transformation as discussed in Ref. [30] to this term reveals that, if the geodesic term is to be regarded as the momentum-space gravitational potential, then the contribution from Eq. (13) can be thought of as arising from a dual drag force in momentum space induced by the scattering. In the next section, we discuss this more within the context of the EFE.
Dissipative Einstein field equations.We now discuss the effect of dissipation on the multiband momentum-space EFE and how it sources the field equations within the present formalism. For convenience, we initially drop band indices from the notation, such that the indices are implied for the various quantities that appear (i.e.,
, etc.). The inverse of the dressed metric is identified as
, such that
, i.e., the dressed (bare) metric raises and lowers indices in the presence (absence) of dissipation. And the screened Levi-Civita connection is defined in the usual sense [38],
, which implies that the dressed Christoffel symbol is that of the dressed metric
This can then be used to derive the dressed Riemann tensor
and its contractions, namely the Ricci tensor
and Ricci scalar
Combining terms, we arrive at the dissipative EFE
where and , which emerge entirely as a result of dissipative scattering, are identified as the local momentum-space cosmological constant and stress tensor, respectively, and are given, up to , by
with the dimension of the system. It should be stressed that this assignment is rather arbitrary as far as the EFE is concerned, and may well be absorbed into the stress tensor [37]. What matters here instead is that the dressed Einstein tensor
no longer vanishes, which physically implies that dissipation is providing a source term for the momentum-space EFE.
To gain further insight into this result, including the quadratic dependence of the source term on the dissipation, it is helpful to recall that the modification of the EFE emerged from the screening of the metric. Following the semiclassical results, one could ask whether there is an entropic explanation for this. To answer this in a consistent manner within the present response-theory formalism, we note that, to second order in the electric field, the rate of entropy production by microscopic scattering processes is determined by the symmetric part of the linear conductivity tensor [55]
which is related to Joule heating and has a quantum geometric contribution that was identified as giving rise to the momentum-space drag force, Eq. (13). Let
denote the increase in entropy as a result of dressing the metric over a time scale given by the scattering time . Restoring band indices and introducing the band resolution of the entropy as
, we find
with
. This demonstrates that the quadratic source term in the multiband EFE can be expressed as a scattering-induced change in entropy associated with interband processes, thereby providing a quantum-response generalization of the arguments in Ref. [30] to dissipative multiband systems.
Conclusion and Outlook.In this work, we have presented an extension of semiclassical momentum-space gravity to dissipative multiband systems by taking a quantum response theory starting point. Within a diagrammatic approach that includes a phenomenological dissipation parameter, we have studied the simultaneous and interconnected dressing of the quantum geometry, carrier dynamics and momentum-space EFE, resulting in a dressed theory of momentum-space gravity. On the technical side, this work introduces and provides two equivalent definitions of the three-state QGT as the simplest quantum geometric quantity beyond the two-state picture, which paves the way for future studies of higher-state quantum geometry. In addition, we have clarified the role of the quantum geometric contorsion tensor and its relation to the full Berry covariant derivative.
Physically, a general viewpoint on quantum response theory that has become more prominent recently is to view it as a probe of the rich Riemannian geometry of quantum state manifolds [31]. The results presented here suggest that this may be harnessed to study a variety of theories of gravitation in quantum materials via optical, magnetic or thermal means once one identifies the corresponding gravitational effect in the multiband system. Another possibly interesting direction to explore from the gravitational point of view is the significance of special disorder scattering (i.e., side jump and skew scattering)–or nonscalar scatterings in the quantum geometry and carrier dynamics [56, 50, 57, 58]. The former is particularly significant, given that special scattering processes in the nonlinear response regime often have relatively complicated and arguably unintuitive classifications. Therefore, finding their gravitational correspondences may be helpful in revealing a more intuitive picture of their nature. Finally, given recent developments in simulating dissipative nonlinear responses via generalizations of quantum phase estimation frameworks [59], it would be quite interesting to consider the possibility of leveraging this approach to simulate theories of gravity via nonlinear responses on quantum computers.
Acknowledgments.We thank Michael Hinczewski, Kurt Hinterbichler and Harsh Mathur for helpful discussions.
Supplemental Material for “Quantum response theory and momentum-space gravity”
S1 S1. Derivation of dressed equation of motion
S1.1 Evaluation of response functions
Using a diagrammatic approach within the Matsubara formalism [51], and applying the Feynman rules presented in Fig. S.1, the linear and quadratic ac conductivities are obtained as
and
where is the frequency of the output photon and are input frequencies. The Matsubara Green’s function is given by
, where is the system Hamiltonian, with
. And for the self-energy, we make the phenomenelogical approximation , with measuring the strength of dissipation induced by scattering events. Finally, velocity operators are obtained by applying successive Berry covariant derivatives,
[…]
Figure S.1: Linear [(a) and (b)] and quadratic [(c)-(f)] conductivity diagrams. Loops and legs are electron and photon propagators, respectively, and vertices imply velocity operator insertions, with the index reserved for the output. The Feynman rules resulting in Eqs. (S.1) and (S.2) consist of the following factors and procedures: 1) A factor of for every set of connected photons. 2) Factors of for the output photon and for each input photon, with . 3) Trace over momentum and band indices of the loop. 4) Matsubara frequency summation .
To evaluate Eqs. (S.1) and (S.2), the effect of the dissipation parameter is included through a phenomenological shift of the frequency factors in the complex plane,
[51]. And the Matsubara Green’s function is expressed in the spectral representation as
where the denominator is the bare Green’s function,
and
is the spectral function [60], with the retarded Green’s function. The frequency summations that are relevant to the evaluation of the conductivities are given by [60, 52]
with the shorthand for the Fermi distribution. In addition, an approximation is needed to obtain analytical results. Taking
results in the identity [52]
where is a general function that includes velocity operator components and the sign is chosen when is analytic in the upper/lower complex plane of the energy integrand. We note that this approximation is not as restricting as may seem, as values as large as have been shown to yield good agreement with numerical results as far as the conductivity is concerned [52]. We apply these steps to Eqs. (9) and (10) and take the dc limit
. After a fair amount of algebra and separating the carrier densities as discussed in the main text, we arrive at the equation of motion, Eq. (11) in the main text.
S1.2 Renormalization functions in the equation of motion
The renormalization functions that appear in Eq. (11) read
while those in Eq. (14) are given by
S2 S2. Derivation of Dissipative field equations
To derive the dissipative field equations, Eq. (17) in the main text, it is helpful to recall the expression for the Levi-Civita connection of an arbitrary rank tensor [61]
as well as the Riemann tensor in terms of connection components
Following the procedure to obtain dressed quantities as discussed in the main text, we perform the necessary contractions to obtain the dressed Ricci tensor and scalar as
and
Combining terms results in the dissipative contributions to the Einstein tensor, , from which the field equations are obtained.